# Study of radon removal performance of silver-ion exchanged zeolite from air for underground experiments

### v1

@@ 1 Introduction
In addition, recent studies have reported that Ag-zeolites have very high adsorption performance with respect to radon in air zeo1 ; zeo2 . It is reported that the radon adsorption performance of Ag-ETS-10 (Engelhard Titano-Silicate, a zeolite analog) and Ag-ZSM-5 (zeolite soconi Mobil-FIve, or Ag-MFI) at room temperature is equivalent to about 500 times that of activated carbon, which could significantly reduce the size of conventional large-scale radon removal systems. Our research group also produced Ag-FER and Ag-MFI in powder form in 2023 and confirmed their ability to adsorb radon in air at room temperature. However, it is also reported in these papers zeo1 ; zeo2 that the adsorption capacity of radon is significantly reduced in high-humidity air. While radon adsorption by granular activated carbon and activated carbon fiber uses physical adsorption due to van-del-Waals forces in the pores, it has been pointed out that the strong adsorption performance of xenon and other noble gases by Ag-zeolite is due to chemical adsorption by silver (metallic silver or silver ions), which has not yet been understood at the molecular level zeo3 . In any case, if chemical adsorption for radon is possible, radon removal should be much more powerful than the traditional radon removal using physical adsorption in the underground experimental field.
@@ 2 Preparation of Ag-FER zeolites
To confirm the distribution of silver inside the pellets, the pellet cross section of silver-exchange zeolite was analyzed by energy dispersive X-ray fluorescence analysis (EDX). Figure 2 shows the results.
Figure 2: EDX measurement results of cross section of silver zeolite pellets. The upper panel shows the distribution of silver and the lower panel shows the distribution of silicon.
@@ 3 Evaluation system for radon adsorption performance
Various measurements were performed using this apparatus under a room temperature environment of about 21–25C. The apparatus was filled with pure air (TAIYO NIPPON SANSO G1 Air) at about atmospheric pressure (differential pressure +0.003 to +0.004 MPaG) as a carrier gas. A mass flow controller (MFC, Horiba SEC-Z500X) and a circulation pump (Enomoto Micro Pump MX-808ST-S) were used to circulate the carrier gas through the system at a flow rate of 1.57 to 3.00 L/min. The temperature of the refrigerator (Taisho TC0147) was set at $-70{}^{\circ}$ to $-90{}^{\circ}$ to control the humidity in the system. Humidity in the carrier gas was monitored with a dew point gauge (Vaisala DMT152). A radon source (PYLON RNC, 78.3 Bq at radiative equilibrium) was connected at all times, and the radon concentration in the system was measured with an 80.3 L radon detector with almost the same performance as the 80 L radon detector developed in Refs. 80Ldet1 ; 80Ldet2 ; Nakano:2017rsy . Ag-zeolites were placed in three nearly identical 1/2-inch stainless steel, 1-m-long U-shaped columns. Figure 5 shows a photograph of the U-shaped column used for the measurement using a 20 g sample.
@@
    If switching to the measurement of a different sample, the U-shaped column was replaced and the Ag-zeolite was baked while evacuating with a turbo molecular pump (TMP). The baking was done at 200C, typically for several hours, and was conducted to an vacuum level of $1\times 10^{-1}$ to $1\times 10^{-3}$ Pa at 200C. Evacuation was continued after the heater was turned off, and finally the vacuum in the column reached $4\sim 8\times 10^{-4}$ Pa at room temperature. The temperature setting of the refrigerator was also changed at this timing if necessary.
@@ 4 Estimation of retention time and radon adsorption coefficient
where $F$ [m3/day] is air flow rate, $RT$ [day] is radon retention time on adsorbent, and $m$ [kg] is mass of adsorbent.
In this study, $RT$ and $K$ were measured and evaluated as follows.
Since the radon sources are always connected, 78.3 Bq of radon is distributed in the system. The radon distribution is assumed to be proportional to the fraction of time when radon stays in each device. The devices with the longest radon stay time in this system are the U-shaped column and the radon detector, and the other devices such as piping are ignored. When the flow rate of the carrier gas is $F$ [m3/day], the time that radon stays in the radon detector ($T$ [day]) is $T=0.0803/F$, given the volume of the radon detector (0.0803 m3). Also, radon should stay in the U-shaped column for the retention time ($RT$ [day]). Therefore, the fraction of radon that stays in the radon detector during circulation through the U-shaped column (called radon concentration ratio: $Rn_{ratio}$) can be expressed as
@@
Therefore, in this study, the measured radon concentration ratio was obtained when the radon concentration was stable, then $RT$ was obtained from Eq. (2), and then $K$ was evaluated from Eq. (1).
@@ 7 Conclusion
As in the preceding studies, radon adsorption in air at room temperature was confirmed. The intrinsic radon emanation from 8Ag-FER-B was evaluated to be less than 45 mBq/kg (90% C.L.). This intrinsic radon emanation level is comparable to those of activated carbons, which have been used in underground experiments. In the case of 8Ag-FER-B, radon removal up to 1 mBq/m3 was observed for about 1 to 2 days after the start of the radon adsorption test, which indicates that 8Ag-FER-B can be applied to underground experiments. On the other hand, radon adsorption efficiency was found to be sensitive with respect to air humidity. In order to increase the efficiency of radon adsorption, dehumidification to a dew point temperature of about $-86{}^{\circ}$ or lower is desirable. To increase the efficiency of radon adsorption, it was effective to increase the amount of silver. Furthermore, it was found that introducing silver in the form of silver ions, rather than metallic silver, is likely to be a good method to increase the efficiency of radon adsorption. The prototype of 8Ag-FER-B in this study showed higher radon adsorption performance than any of the preceding studies, and when 8Ag-FER-B is used in the Hyper-Kamiokande experiment, it is expected to be much smaller (12-24 kg) than the existing radon removal system using room-temperature activated carbon.
@@ Appendix A Measured data
Figure 9: 8Ag-FER-B 44g, 1.57 SLM, $-90{}^{\circ}$ (type and amount of Ag-zeolite, carrier gas flow rate, refrigerator temperature) measurement. Bottom panel: Variation of the radon concentration. Horizontal axis is elapsed day from the U-shaped column connection. The left (right) vertical axis is radon concentration in the radon detector (Rn adsorption coefficient). The red line shows the fitting result with the Rn source radiation equilibrium function. The blue line shows the extrapolation of the red line. Top panel: Variation of the dew point temperature. The horizontal axis is common with the bottom plot. The hatched area was not available to acquire data due to a planned power cut.

### v2

@@ 1 Introduction
In addition, recent studies have reported that Ag-zeolites have very high adsorption performance with respect to radon in air zeo1 ; zeo2 . It is reported that the radon adsorption performance of Ag-ETS-10 (Engelhard Titano-Silicate, a zeolite analog) and Ag-ZSM-5 (zeolite soconi Mobil-FIve, or Ag-MFI) at room temperature is equivalent to about 500 times that of activated carbon, which could significantly reduce the size of conventional large-scale radon removal systems. Our research group also produced Ag-FER and Ag-MFI in powder form in 2023 and confirmed their ability to adsorb radon in air at room temperature ogawa2 . However, it is also reported in these papers zeo1 ; zeo2 that the adsorption capacity of radon is significantly reduced in high-humidity air. While radon adsorption by granular activated carbon and activated carbon fiber uses physical adsorption due to van-del-Waals forces in the pores, it has been pointed out that the strong adsorption performance of xenon and other noble gases by Ag-zeolite is due to chemical adsorption by silver (metallic silver or silver ions), which has not yet been understood at the molecular level zeo3 . In any case, if chemical adsorption for radon is possible, radon removal should be much more powerful than the traditional radon removal using physical adsorption in the underground experimental field.
@@ 2 Preparation of Ag-FER zeolites
To confirm the distribution of silver inside the pellets, the pellet cross section of Ag-zeolite was analyzed by energy dispersive X-ray fluorescence analysis (EDX). Figure 2 shows the results.
Figure 2: EDX measurement results of cross section of Ag-zeolite pellets. The upper panel shows the distribution of silver and the lower panel shows the distribution of silicon.
@@ 3 Evaluation system for radon adsorption performance
Various measurements were performed using this apparatus under a room temperature environment of about 21–25C. The apparatus was filled with pure air (TAIYO NIPPON SANSO G1 Air) at about atmospheric pressure (differential pressure +0.003 to +0.004 MPaG) as a carrier gas. A mass flow controller (MFC, Horiba SEC-Z500X) and a circulation pump (Enomoto Micro Pump MX-808ST-S) were used to circulate the carrier gas through the system at a flow rate of 1.57 to 3.00 L/min. The temperature of the refrigerator (Taisho TC0147) was set at $-70{}^{\circ}$ to $-90{}^{\circ}$ to control the humidity in the system. Humidity in the carrier gas was monitored with a dew point gauge (Vaisala DMT152). A radon source (PYLON RNC, 78.3 Bq at radiative equilibrium) was connected at all times, and the radon concentration in the system was measured with an 80.3 L radon detector with almost the same performance as the 80 L radon detector developed in Refs. 80Ldet1 ; 80Ldet2 ; Nakano:2017rsy . Ag-zeolites were placed in three nearly identical 1/2-inch stainless steel, 1-m-long U-shaped columns. Figure 5 shows a photograph of the U-shaped columns used for the measurement using a 20 g sample.
@@
    If switching to the measurement of a different sample, the U-shaped column was replaced and the Ag-zeolite was baked while evacuating with a turbo molecular pump (TMP). The baking temperature was set at 200C, typically for several hours, and was conducted to an vacuum level of $1\times 10^{-1}$ to $1\times 10^{-3}$ Pa at 200C. Evacuation was continued after the heater was turned off, and finally the vacuum in the column reached $4\times 10^{-4}\sim 8\times 10^{-4}$ Pa at room temperature. The temperature setting of the refrigerator was also changed at this timing if necessary.
@@ 4 Estimation of retention time and radon adsorption coefficient
where $F$ [m3/day] is air flow rate, $RT$ [day] is radon retention time on adsorbent, and $m$ [kg] is mass of adsorbent. In this study, $RT$ and $K$ were measured and evaluated as follows.
Since the radon source is always connected, 78.3 Bq of radon is distributed in the system when radiative equilibrium. The radon distribution is assumed to be proportional to the fraction of time when radon stays in each device [^1] Past measurements have confirmed that this assumption is valid to some extent.. The devices with the longest radon stay time in this system are the U-shaped column and the radon detector, and the other devices such as piping are ignored. When the flow rate of the carrier gas is $F$ [m3/day], the time that radon stays in the radon detector ($T$ [day]) is $T=0.0803/F$, given the volume of the radon detector (0.0803 m3). Also, radon should stay in the U-shaped column for the retention time ($RT$ [day]). Therefore, the fraction of radon that stays in the radon detector during circulation through the U-shaped column (called radon concentration ratio: $Rn_{ratio}$) can be expressed as
@@
When calculating $Rn_{ratio}$ from a measurement, the radiative equilibrium curve is used as the denominator in Eq. (3). Therefore, in this study, the measured radon concentration ratio was obtained when the radon concentration was stable, then $RT$ was obtained from Eq. (2), and then $K$ was evaluated from Eq. (1).
@@ 7 Conclusion
As in the preceding studies, radon adsorption in air at room temperature was confirmed. The intrinsic radon emanation from 8Ag-FER-B was evaluated to be less than 45 mBq/kg (90% C.L.). This intrinsic radon emanation level is comparable to those of activated carbons, which have been used in underground experiments. In the case of 8Ag-FER-B, radon removal down to 1 mBq/m3 was observed for about 1 to 2 days after the start of the radon adsorption test, which indicates that 8Ag-FER-B can be applied to underground experiments. On the other hand, radon adsorption efficiency was found to be sensitive with respect to air humidity. In order to increase the efficiency of radon adsorption, dehumidification to a dew point temperature of about $-86{}^{\circ}$ or lower is desirable. To increase the efficiency of radon adsorption, it was effective to increase the amount of silver. Furthermore, it was found that introducing silver in the form of silver ions, rather than metallic silver, is likely to be a good method to increase the efficiency of radon adsorption. The prototype of 8Ag-FER-B in this study showed higher radon adsorption performance than any of the preceding studies, and when 8Ag-FER-B is used in the Hyper-Kamiokande experiment, it is expected to be much smaller (12-24 kg, 17-35 L in volume) than the existing radon removal system using room-temperature activated carbon.
@@ Appendix A Measured data
Figure 9: 8Ag-FER-B 44g, 1.57 SLM, $-90{}^{\circ}$ (type and amount of Ag-zeolite, carrier gas flow rate, refrigerator temperature) measurement. Bottom panel: Variation of the radon concentration. Horizontal axis is elapsed day from the U-shaped column connection. The left (right) vertical axis is radon concentration in the radon detector (Rn adsorption coefficient). The red line shows the fitting result with the Rn source radiative equilibrium function. The blue line shows the extrapolation of the red line. Top panel: Variation of the dew point temperature. The horizontal axis is common with the bottom plot. The hatched area was not available to acquire data due to a planned power cut.